OpenStudy (anonymous):

As a summer job, Randall Jui can work at the amusement park earning $8.75 an hour for 20 hours a week, or as a busboy in a restaurant. As a busboy, he would earn $8.00 per hour for the same 20-hour week. How much more would Randall earn at the amusement park than at the restaurant in one week?

OpenStudy (texaschic101):

8.75 * 20 = $ 175 working at amusement park in 1 week 8.00 * 20 = $ 160 working as a busboy in 1 week 175 - 160 = $ 15 more working at amusement park